# Evaluating The Explainability of State-of-the-Art Deep Learning-based Network Intrusion Detection Systems

### v3

@@ VI Discussion and Future Work
From the experimental results obtained in the previous section, we have seen that some DL-based NIDS models (Kitsune, ENIDrift) can be better interpreted in terms of decision trees than other models (HorusEye, HELAD). This might be attributed to the complexity of the NIDS model architecture and the properties of the model components. For example, HorusEye NIDS fidelity values are in the negative. Though HorusEye architecture includes only a single autoencoder in the second stage (which processes control plane traffic) compared to an ensemble of AEs in Kitsune or ENIDrift, that single AE is modified from a conventional AE to keep it lightweight, reduce the computational complexity, reduce the false-positive rate and improve throughput. It uses an asymmetric AE with separable convolution, dilation convolution and model quantization. As it is, AEs are limited in their interpretability as a result of their non-linear nature. Due to the complex architecture of the modified AE deployed in HorusEye, TRUSTEE might not be able to build a DT which can imitate the decisions of the black-box modified AE within an acceptable error rate. Therefore, to build DL-based NIDSs with better “explainability”, we recommend using less complex DL architectures and conventional components. However, there will be a trade-off between “explainability” and model performance (e.g., false-positive rate) and NIDS designers need to achieve a balance between the two.
@@
Finally, it can be inferred from the results that some DL-based NIDSs (Kitsune) are more vulnerable to variable packet rate samples than others (HorusEye, HELAD) while for the rest (ENIDrift), such vulnerability remains inconclusive. One of the possible underlying reasons is that sometimes NIDSs are trained with training data containing higher than real-world attack traffic volume. To reduce vulnerability to variable packet rate samples, we recommend that NIDS designers use training data which reflects the attack traffic volume in real-world deployments. Additionally, the NIDS architecture may also have a role to play in its vulnerability to variable packet rate samples. An AE trained on benign traffic may be vulnerable to attack traffic with variable packet rates as the threshold used for output loss function (RMSE) is determined based on “in-distribution” attack traffic present in the training data. Since, Kitsune consists of an ensemble of AEs, the vulnerability effect is amplified. HorusEye consists of a single modified AE and therefore, its vulnerability to variable packet rate samples is not as pronounced as Kitsune. ENIDrift uses a different method of feature extraction (vector embedding) and updates the weights assigned to AEs in the ensemble based on change in attack traffic distribution which might keep it from producing a discernable pattern of dependence on attack traffic volume compared to Kitsune. We have observed earlier that HELAD NIDS anomaly scores are not vulnerable to variable packet rate samples to a significant extent. This may be attributed to HELAD’s architecture which combines an AE with an LSTM network. Since the final anomaly score is a weighted exponential aggregation of AE’s actual RMSE score and LSTM’s predicted RMSE scores for future samples (using historical samples). Even if the AE is vulnerable to variable packet rate samples, the LSTM’s predicted RMSE scores might be much larger compared to to the AE’s RMSE score (the exponential operation makes the value even larger), thus affecting the final anomaly score to a greater extent and reducing HELAD’s vulnerability to variable packet rate samples.
VII Conclusion
@@
We have analyzed four state-of-the-art DL-based NIDS models (Kitsune, HorusEye, ENIDRift, HELAD) using explainable AI techniques such as TRUSTEE and SHAP. Using both global explanations and local explanations for the blackbox models’ decisions, we have presented the most prominent features used by each NIDS model considered. We have compared the explanations generated across xAI methods for a given NIDS model (using our devised data subset-level explanation approach) as well as the explanations generated across the NIDS models for a given xAI method (using fidelity and other standard evaluation metrics). Finally, we evaluate the vulnerability of each NIDS model to inductive bias (variable packet rate samples). The results show that: (1) some DL-based NIDS models can be better interpreted in terms of decision trees than other models, (2) TRUSTEE and SHAP explanations are in conflict for most of the NIDS models considered in this work and (3) some NIDS models are more vulnerable to variable packet rate samples than other models.
@@ Appendix A More Experimental Results
A-A Case Study- Kitsune $\times$ CICIDS-2017
TRUSTEE Analysis: The results of TRUSTEE analysis for Kitsune are shown in Table VII. Using TRUSTEE with $30\%$ of the original CICIDS-2017 dataset samples and no pruning results in the DT explanation which achieves 0.843 fidelity compared to Kitsune. Using TRUSTEE with $30\%$ of the dataset samples and top-k Pruning method (setting $k=3$) results in a DT explanation that is shown in Fig. 13 and achieves 0.919 fidelity compared to Kitsune.
The top three prominent features that Kitsune uses to determine an anomaly are: $HpHp\_0.1\_pcc\_0\_1$, $HH\_0.1\_pcc\_0\_1$ and $HH\_jit\_0.01\_std$. Thus, Kitsune relies mainly on the sizes of packets and the jitter experienced by packets exchanged per time frame to determine if an attack is underway.
Sample size Top-k pruning used? DT size, depth, leaves Fidelity
@@ A-A Case Study- Kitsune $\times$ CICIDS-2017
TABLE VII: TRUSTEE analysis results for Kitsune CICIDS-2017 model
@@
SHAP Analysis: Using kernel SHAP analysis, Fig. 14 shows a summary of how the top features in CICIDS-2017 dataset impact the Kitsune NIDS model’s output through a beeswarm plot for randomly selected 1000 data samples corresponding to the benign root node-leaf node path in the TRUSTEE DT (Fig. 13). Features such as $HpHp\_0.01\_weight\_0$, $HH\_jit\_0.01\_weight$ and $HH\_0.01\_weight\_0$ have the most impact on the prediction value, in that order.
TRUSTEE vs SHAP Analysis: Using the systematic approach for comparing TRUSTEE and SHAP explanations as explained in sub-section IV-B, since the set of features used in the benign root node-leaf node path in the TRUSTEE DT is not a subset of the set of top contributing features computed by kernel SHAP summary analysis ($HpHp\_0.1\_pcc\_0\_1$ is the only common feature between the two sets), TRUSTEE AND SHAP explanations do not agree with each other at the subset level for Kitsune NIDS.
@@
Detection of Inductive Bias: The CICIDS-2017 dataset does not provide a separate dataset for botnet traffic trace which could be used for comparison with the Mirai trace in terms of the attack traffic volume. However, using a similar approach for tampering as in the case of Mirai trace, we evaluated the following threshold ranges (in packets per second) for ARP requests originating from infected hosts : (1) 10-50, (2) 30-70 and (3) 50-90. The original trace as well as the tampered traces are shown in Fig. 15. It can be observed that: (1) there is no significant difference between the benign and attack traffic volumes for the original trace and (2) tampering had negligible effect on the difference between benign and attack traffic in terms of number of packets per second. Thus, unlike the Mirai trace, there is no significant difference between the benign and attack traffic volumes which can be learnt by the NIDS model and leave it vulnerable to a change in attack traffic distribution. Thus, we do not investigate the presence of inductive bias (vulnerability to variable packet rate samples) any further for NIDS models trained with the CICIDS-2017 dataset.
Figure 15: Number of packets per second for original CICIDS-2017 trace and tampered traces. Blue portions of the plot correspond to benign traffic while red portions correspond to malicious traffic inclusive of benign and attack traffic.
A-B Case Study- HorusEye $\times$ CICIDS-2017
TRUSTEE Analysis: The results of TRUSTEE analysis for HorusEye are shown in Table VIII. Using TRUSTEE with $30\%$ of the CICIDS-2017 dataset samples with no pruning results in a DT explanation which achieves negative fidelity ($\sim-0.956$) compared to original HorusEye model. Using TRUSTEE with $30\%$ of the dataset samples and top-k Pruning method (setting $k=10$) results in a DT explanation which achieves negative fidelity ($\sim-0.018$) compared to original HorusEye model but higher fidelity compared to no pruning. Negative fidelity values mean that the DT explanation fits HorusEye predictions worse than the mean of the predictions.
Nevertheless, the decision tree for HorusEye Mirai model with top-k pruning (k=10) is shown in Fig. 16. The top three prominent features that HorusEye’s DT uses to determine an anomaly are: $HH\_5\_pcc\_0\_1$, $HpHp\_1\_covariance\_0\_1$ and $HH\_jit\_0.01\_std$. Thus, HorusEye’s DT relies mainly on the sizes of packets and the jitter experienced by packets exchanged between a pair of IP addresses per time frame to determine if an attack is underway.
Sample size Top-k pruning used? DT size, depth, leaves Fidelity
@@ A-B Case Study- HorusEye $\times$ CICIDS-2017
TABLE VIII: TRUSTEE analysis results for HorusEye CICIDS-2017 model
@@
SHAP Analysis: Using kernel SHAP analysis, Fig. 17 shows a summary of how the top features in CICIDS-2017 dataset impact the HorusEye NIDS model’s output using a beeswarm plot for randomly selected 1000 data samples corresponding to the benign root node-leaf node path in the TRUSTEE DT (Fig. 16). Features such as $HH\_jit\_0.1\_std$, $MI\_dir\_0.01\_std$ and $HpHp\_0.1\_weight\_0$ have the most impact on the prediction value, in that order.
@@
TRUSTEE vs SHAP Analysis: Using the systematic approach for comparing TRUSTEE and SHAP explanations as explained in sub-section IV-B, the set of features used in the benign root node-leaf node path in the TRUSTEE DT is not a subset of the set of top contributing features computed by kernel SHAP summary analysis. Therefore, TRUSTEE AND SHAP explanations do not agree with each other at the subset level for HorusEye NIDS.
A-C Case Study- ENIDrift $\times$ CICIDS-2017
@@
TRUSTEE Analysis: The results of TRUSTEE analysis for ENIDrift are shown in Table IX. Using TRUSTEE with $30\%$ of the CICIDS-2017 dataset samples and no pruning results in the DT explanation which achieves 0.929 fidelity compared to ENIDrift. Using TRUSTEE with $30\%$ of the dataset samples and top-k Pruning method (setting $k=10$) results in a DT explanation (shown in Fig. 18) which achieves 0.53 fidelity compared to ENIDrift. This is expected since pruning the branches of a DT reduces its fidelity. The most prominent features ENIDrift uses to determine anomaly are: $f38$, $f81$, $f158$.
Sample size Top-k pruning used? DT size, depth, leaves Fidelity
@@
TABLE IX: TRUSTEE analysis results for ENIDrift CICIDS-2017 model
@@
TRUSTEE vs SHAP Analysis: Using the systematic approach for comparing TRUSTEE and SHAP explanations as explained in sub-section IV-B, the set of features used in the benign root node-leaf node path in the TRUSTEE DT is not a subset of the set of top contributing features computed by kernel SHAP summary analysis (there are no common features between the two sets). Therefore, TRUSTEE AND SHAP explanations do not agree with each other at the subset level for ENIDrift NIDS.
A-D Case Study- HELAD $\times$ CICIDS-2017
@@
TRUSTEE Analysis: The results of TRUSTEE analysis for HELAD are shown in Table X. Using TRUSTEE with $30\%$ of the reduced CICIDS-2017 dataset samples and no pruning results in the DT explanation which achieves 0.0 fidelity compared to HELAD. Using TRUSTEE with $30\%$ of the dataset samples and top-k Pruning method (setting $k=10$) results in a DT explanation (shown in Fig. 20) which achieves 0.0 fidelity compared to HELAD.
@@
    $HH\_jit\_0.01\_mean$- Packet jitter aggregated by traffic sent between a set of source and destination IP addresses with time window 0.01 (1 minute).
@@
    $HH\_0.1\_radius\_0\_1$- L2-norm of variances of two packet size streams aggregated by traffic sent between a set of source and destination IP addresses with time window 0.1 (10 seconds).
@@
    $MI\_dir\_1.0\_std$- Standard deviation of packet sizes aggregated by source MAC and IP addresses, with time window 1.0 (1.5 seconds).
@@
Sample size Top-k pruning used? DT size, depth, leaves Fidelity
@@
TABLE X: TRUSTEE analysis results for HELAD NIDS

### v4

@@ 5 Discussion and Future Work
From the experimental results obtained in the previous section, we have seen that some DL-based NIDS models (Kitsune, ENIDrift) can be better interpreted in terms of decision trees than other models (HorusEye, HELAD). This might be attributed to the complexity of the NIDS model architecture and the properties of the model components. For example, HorusEye NIDS fidelity values are in the negative. Though HorusEye architecture includes only a single autoencoder in the second stage (which processes control plane traffic) compared to an ensemble of AEs in Kitsune or ENIDrift, that single AE is modified from a conventional AE to keep it lightweight, reduce the computational complexity, reduce the false-positive rate and improve throughput. It uses an asymmetric AE with separable convolution, dilation convolution and model quantization. As it is, AEs are limited in their interpretability as a result of their non-linear nature. Due to the complex architecture of the modified AE deployed in HorusEye, TRUSTEE might not be able to build a DT which can imitate the decisions of the black-box modified AE within an acceptable error rate. Therefore, to build DL-based NIDSs with better “explainability”, we recommend using less complex DL architectures and conventional components. However, there will be a trade-off between “explainability” and model performance (e.g., false-positive rate) and NIDS designers need to achieve a balance between the two. One of the ways to achieve this balance is to use the weighted metric introduced in Section 4.2.
@@
Zhao et al. [32] have shown that explanations can leak information about the black-box model. This information can be used for model inversion attacks which reconstruct input data using corresponding predictions and model explanation. The robustness of explanation techniques to model inversion attacks is an important criterion, especially when it comes to NIDSs. This is because if an NIDS model can be reconstructed, it can be used by attackers to craft attack traffic to evade detection by that NIDS. However, there exist techniques [33], limited as they may be, to generate inversion-resistant model explanations. Again, it should be noted that all of the above works are focused on image classification. Further work is needed to assess the robustness of explanation methods such as SHAP and TRUSTEE to being exploited for black-box model inversion.
6 Related Work
As a result, we can see the beginning of efforts in the research community to apply XAI techniques to ML-based security systems. A few works [9, 10, 11, 12] have employed XAI methods such as LIME [14] and SHAP [8] to assess the quality of local/global explanations generated for conventional machine learning classifiers (AdaBoost, k-nearest neighbour, Multi-layer Perceptron, Random Forest, XGBoost, Light Gradient-Boosting Machine, Gradient Boosting Classifier) trained on network intrusion datasets (NSL-KDD dataset [36], CICIDS-2017 [28]) in terms of criteria such as faithfulness, stability, complexity, and sensitivity. In [13], the authors have delved into six explanation methods, evaluating their effectiveness across four security systems described in existing literature. These systems leverage DL techniques to detect Android malware, malicious PDF files and security vulnerabilities. The assessment criteria encompass both general properties of deep learning and domain-specific aspects relevant to security.
  • Existing works that evaluate XAI methods on ML/DL-based NIDSs ([9, 10, 11, 12, 38, 40]): (1) do not compare the explainability of those NIDSs with that of state-of-the-art DL-based NIDSs and (2) do not include the latest XAI tools for ML/DL-based network security such as TRUSTEE [7].
7 Conclusion
@@
We have analyzed four state-of-the-art DL-based NIDS models (Kitsune, HorusEye, ENIDRift, HELAD) using XAI techniques such as TRUSTEE and SHAP. We have compared the explanations generated across XAI methods using our proposed TRUSTEE-SHAP agreement score as well as other security-focused criteria. Using both global explanations and local explanations for the blackbox models’ decisions, we have presented the most prominent features used by each NIDS model considered. The results show that: (1) some DL-based NIDS models can be better interpreted in terms of decision trees than other models, (2) TRUSTEE and SHAP explanations are in conflict for most of the NIDS models considered in this work and (3) both TRUSTEE and SHAP explanations are vulnerable to adversarial attacks though TRUSTEE provides better stability than SHAP.
In the future, we plan to design an NIDS using DL architectures of lower complexity and compare its “explainability” with other state-of-the-art NIDSs using appropriate metrics. We also plan to investigate the reason(s) behind the lack of agreement between TRUSTEE and SHAP explanations for an NIDS at the data subset-level and use that information to improve the level of agreement.
@@ Appendix A More Experimental Results
A.1 Case Study- Kitsune $\times$ CICIDS-2017
TRUSTEE Analysis: The results of TRUSTEE analysis for Kitsune are shown in Table 8. Using TRUSTEE with $30\%$ of the original CICIDS-2017 dataset samples and no pruning results in the DT explanation which achieves 0.843 fidelity compared to Kitsune. Using TRUSTEE with $30\%$ of the dataset samples and top-k Pruning method (setting $k=3$) results in a DT explanation that is shown in Fig. 9 and achieves 0.919 fidelity compared to Kitsune.
The top three prominent features that Kitsune uses to determine an anomaly are: HpHp_0.1_pcc_0_1HpHp\_0.1\_pcc\_0\_1, HH_0.1_pcc_0_1HH\_0.1\_pcc\_0\_1 and HH_jit_0.01_stdHH\_jit\_0.01\_std. Thus, Kitsune relies mainly on the sizes of packets and the jitter experienced by packets exchanged per time frame to determine if an attack is underway.
Sample size
Top-k pruning
used?
DT size, depth, leaves Fidelity
@@ A.1 Case Study- Kitsune $\times$ CICIDS-2017
Table 8: TRUSTEE analysis results for Kitsune CICIDS-2017 model
@@
SHAP Analysis: Using kernel SHAP analysis, Fig. 10 shows a summary of how the top features in CICIDS-2017 dataset impact the Kitsune NIDS model’s output through a beeswarm plot for randomly selected 1000 data samples corresponding to the benign root node-leaf node path in the TRUSTEE DT (Fig. 9). Features such as HpHp_0.01_weight_0HpHp\_0.01\_weight\_0, HH_jit_0.01_weightHH\_jit\_0.01\_weight and HH_0.01_weight_0HH\_0.01\_weight\_0 have the most impact on the prediction value, in that order.
@@
A.2 Case Study- HorusEye $\times$ CICIDS-2017
TRUSTEE Analysis: The results of TRUSTEE analysis for HorusEye are shown in Table 9. Using TRUSTEE with $30\%$ of the CICIDS-2017 dataset samples with no pruning results in a DT explanation which achieves negative fidelity ($\sim-0.956$) compared to original HorusEye model. Using TRUSTEE with $30\%$ of the dataset samples and top-k Pruning method (setting $k=10$) results in a DT explanation which achieves negative fidelity ($\sim-0.018$) compared to original HorusEye model but higher fidelity compared to no pruning. Negative fidelity values mean that the DT explanation fits HorusEye predictions worse than the mean of the predictions.
Nevertheless, the decision tree for HorusEye Mirai model with top-k pruning (k=10) is shown in Fig. 11. The top three prominent features that HorusEye’s DT uses to determine an anomaly are: HH_5_pcc_0_1HH\_5\_pcc\_0\_1, HpHp_1_covariance_0_1HpHp\_1\_covariance\_0\_1 and HH_jit_0.01_stdHH\_jit\_0.01\_std. Thus, HorusEye’s DT relies mainly on the sizes of packets and the jitter experienced by packets exchanged between a pair of IP addresses per time frame to determine if an attack is underway.
Sample size
Top-k pruning
used?
DT size, depth, leaves Fidelity
@@ A.2 Case Study- HorusEye $\times$ CICIDS-2017
Table 9: TRUSTEE analysis results for HorusEye CICIDS-2017 model
@@
SHAP Analysis: Using kernel SHAP analysis, Fig. 12 shows a summary of how the top features in CICIDS-2017 dataset impact the HorusEye NIDS model’s output using a beeswarm plot for randomly selected 1000 data samples corresponding to the benign root node-leaf node path in the TRUSTEE DT (Fig. 11). Features such as HH_jit_0.1_stdHH\_jit\_0.1\_std, MI_dir_0.01_stdMI\_dir\_0.01\_std and HpHp_0.1_weight_0HpHp\_0.1\_weight\_0 have the most impact on the prediction value, in that order.
@@
A.3 Case Study- ENIDrift $\times$ CICIDS-2017
@@
TRUSTEE Analysis: The results of TRUSTEE analysis for ENIDrift are shown in Table 10. Using TRUSTEE with $30\%$ of the CICIDS-2017 dataset samples and no pruning results in the DT explanation which achieves 0.929 fidelity compared to ENIDrift. Using TRUSTEE with $30\%$ of the dataset samples and top-k Pruning method (setting $k=10$) results in a DT explanation (shown in Fig. 13) which achieves 0.53 fidelity compared to ENIDrift. This is expected since pruning the branches of a DT reduces its fidelity. The most prominent features ENIDrift uses to determine anomaly are: $f38$, $f81$, $f158$.
Sample size
Top-k pruning
used?
DT size, depth, leaves Fidelity
@@
Table 10: TRUSTEE analysis results for ENIDrift CICIDS-2017 model
@@
A.4 Case Study- HELAD $\times$ CICIDS-2017
@@
TRUSTEE Analysis: The results of TRUSTEE analysis for HELAD are shown in Table 11. Using TRUSTEE with $30\%$ of the reduced CICIDS-2017 dataset samples and no pruning results in the DT explanation which achieves 0.0 fidelity compared to HELAD. Using TRUSTEE with $30\%$ of the dataset samples and top-k Pruning method (setting $k=10$) results in a DT explanation (shown in Fig. 15) which achieves 0.0 fidelity compared to HELAD.
@@
    HH_jit_0.01_meanHH\_jit\_0.01\_mean- Packet jitter aggregated by traffic sent between a set of source and destination IP addresses with time window 0.01 (1 minute).
@@
    HH_0.1_radius_0_1HH\_0.1\_radius\_0\_1- L2-norm of variances of two packet size streams aggregated by traffic sent between a set of source and destination IP addresses with time window 0.1 (10 seconds).
@@
    MI_dir_1.0_stdMI\_dir\_1.0\_std- Standard deviation of packet sizes aggregated by source MAC and IP addresses, with time window 1.0 (1.5 seconds).
@@
Sample size
Top-k pruning
used?
DT size, depth, leaves Fidelity
@@
Table 11: TRUSTEE analysis results for HELAD NIDS